A light bulb company produces a constant number of new light bulbs in their factory each week. The company stores the bulbs in a warehouse where old light bulbs from the previous year are also stored. After 2 weeks, the company has 5,000 bulbs in the warehouse. After 7 weeks, the company has 8,500 bulbs in the warehouse.

A: What is the constant rate of change in this problem?
B: Which equations correctly model the number of bulbs in the warehouse as a function of the number of weeks?
One answer choice for question A, and select all correct answer choices for question B.
B: y+8500=700(x+7)
A: 700 bulbs per week
B: y+5000=700(x+2)
B: y−8500=−700(x−7)
A: −700 weeks per bulb
B: y+5000=−700(x+2)
B: y−5000=−700(x−2)
A: 700 weeks per bulb
B: y−5000=700(x−2)
A: −700 bulbs per week
B: y+7=−700(x+8500)
B: y−8500=700(x−7)

Answer by stanbon(75887)

A light bulb company produces a constant number of new light bulbs in their factory each week. The company stores the bulbs in a warehouse where old light bulbs from the previous year are also stored. After 2 weeks, the company has 5,000 bulbs in the warehouse. After 7 weeks, the company has 8,500 bulbs in the warehouse.
-----
Note: You have 2 points:: (2,5000) and (7,8500)
----------------------
A: What is the constant rate of change in this problem?
slope = (8500-5000)/(7-2) = 3500/5 = 700
-----------------
B: Which equations correctly model the number of bulbs in the warehouse as a function of the number of weeks?
y-8500 = 700(x-7)
y-5000 = 700(x-2)
------------
Cheers,
Stan H.
------
One answer choice for question A, and select all correct answer choices for question B.
B: y+8500=700(x+7)
A: 700 bulbs per week
B: y+5000=700(x+2)
B: y−8500=−700(x−7)
A: −700 weeks per bulb
B: y+5000=−700(x+2)
B: y−5000=−700(x−2)
A: 700 weeks per bulb
B: y−5000=700(x−2)
A: −700 bulbs per week
B: y+7=−700(x+8500)
B: y−8500=700(x−7)